OpenStudy (anonymous):

can anyone help me with this question or not y/z divide z/t = 1 for y

OpenStudy (anonymous):

y/z / z/t = 1 y/z into t/z = 1 yt/z^2 = 1 y=z^2/t

OpenStudy (anonymous):

thanks